OpenStudy (anonymous):

Which of the following is the additive inverse of -4b? 4b -4b<----my answer 1/4b -1/4b Which values for x and y make the statement (x + 5)(y - 6) = 0 true? x = 5, y = -6 x = 5, y = 6 x = -5, y = 6 x = -5, y = -6

hartnn (hartnn):

additive inverse of x is -x that is we just switch the signs so additive inverse of -4b will be??

OpenStudy (anonymous):

opps i meant i chose A my mistake

hartnn (hartnn):

when a+b = 0 we say a=-b when pq = 0, we say either p=0 or q=0 so, when (x+5)(y-6) =0 we say x+5 = 0 or y-6 = 0 x+5 = 0 just subtract 5 from both sides, what do u get? y-6 =0 just add 6 on both sides, what do u get?

OpenStudy (anonymous):

so x= -5 and y= 6?

hartnn (hartnn):

yes, 4b is correct for the 1st one and x=-5, y= 6 is also correct! good :0

OpenStudy (anonymous):

thank you cx

OpenStudy (anonymous):

Which values for a and b make the statement (a - 2)(b + 3) = 0 true? a = -2, b = 3 a = 2, b = 3 a = -2, b = -3 a = 2, b = -3

OpenStudy (anonymous):

is the answer to this question D?

hartnn (hartnn):

yes, D is correct :)

OpenStudy (anonymous):

ok thank you again cx